OpenStudy (anonymous):

if a and b are the roots of the equation 2x^2 - px + 7 =0 , a/b is a root of the equation (a) 7x^2 + 28x - p^2x + 14 =0 (b) 14x^2 + (28-p^2)x + 14 =0

OpenStudy (anonymous):

from what the roots are we can infer that\[(2x-2a)(x-b)=0\]from multiplying it out and combining like terms we know that\[-2b-2a=-p\]and the final coefficient is\[2ab=7\]

OpenStudy (anonymous):

\[2a+2b=p\]\[ab=\frac72\]

OpenStudy (anonymous):

we can eliminate at least one of the variables at this point

OpenStudy (anonymous):

\[ab=7\implies b=\frac7{2a}\]put this into the first equation\[2a+\frac7a=p\]\[2a^2+7a-pa=0\]\[a[2a+7-p]=0\]hm... still not drawing any conclusions

OpenStudy (anonymous):

\(a\) can't be zero from the first line in the above post, so that leaves that\[2a+7-p=0\]\[a={7-p\over2}\]we could have done the same argument with b, so that seems to mean that\[a=b=\frac{7-p}2\]

OpenStudy (anonymous):

somebody feel free to criticize what I'm doing or jump in or whatever

OpenStudy (asnaseer):

you got:\[2a+2b=p\]and:\[2ab=7\]therefore:\[2(a+b)=p\implies4(a+b)^2=p^2\implies4(a^2+b^2+2ab)=p^2\]therefore:\[4(a^2+b^2+7)=p^2\implies a^2+b^2=\frac{p^2}{4}-7\]

OpenStudy (asnaseer):

now, lets say that the other quadratic has a roots of a/b and say r, and that it is of the form:\[x^2-ux+v=0\]then this implies:\[\frac{ar}{b}=u\]and\[\frac{a}{b}+r=v\]

OpenStudy (asnaseer):

this leads to:\[r=\frac{ub}{a}\]

OpenStudy (asnaseer):

therefore:\[\frac{a}{b}+\frac{ub}{a}=v\]multiplying both sides by ab gives:\[a^2+ub^2=vab\]

OpenStudy (asnaseer):

the first equation:\[14x^2 + (28-p^2)x + 14 =0\]can be divided by 14 to give:\[x^2-(\frac{p^2}{14}-2)x+1=0\]which gives:\[v=\frac{p^2}{14}\]and:\[u=1\]

OpenStudy (asnaseer):

sorry I had u and v the wrong way around up there

OpenStudy (experimentx):

oh .. simply replace 'p' by 'a, b' in those options ... and try to factor out (x-a/b) from those options

OpenStudy (asnaseer):

I should have initially said the equation is:\[x^2-vx+u=0\]

OpenStudy (asnaseer):

this then gives:\[a^2+ub^2=vab\]leading to:\[a^2+b^2=(\frac{p^2}{14}-2)ab=(\frac{p^2}{14}-2)\times\frac{7}{2}=\frac{p^2}{4}-7\]

OpenStudy (asnaseer):

which meets the original equation we had for \(a^2+b^2\)

OpenStudy (asnaseer):

I made /some/ algebra mistakes up there but hopefully it all adds up correctly

OpenStudy (asnaseer):

@Yahoo! did you follow that or do you want to explain anything more?

OpenStudy (asnaseer):

*want me to...

OpenStudy (experimentx):

you have  right??